OpenStudy (anonymous):

solve the equation. 1. x-6=8 a. x=-2 b. x=2 c. x=-14 d. x=14

OpenStudy (anonymous):

x-6=8 +6 +6 x=14

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

no problem

OpenStudy (anonymous):

more simply x-6=8 x=8+6 x=14